A chemist wants to obtain 80 liters of a solution is 12% hydrochloric acid. How many liters of 10% hydrochloric acid solution and 20% hydrochloric acid solution should he mix to do so

Answer by checkley79(3341)

.10X+.20(80-X)=.12*80
.10X+16-.20X=9.6
-.10X=9.6-16
-.10X=-6.4
X=-6.4/-.10
X=64 LITERS OF 10% ACID IS USED.
80-64=16 LITERS OF 20% IS USED.
PROOF:
.10*64+.20*16=9.6
6.4+3.2=9.6
9.6=9.6